gina has $2 more than three times the amount fred has. when gina gave fred $5 fred then had one half asa much as gina. how much did gina have at first
Answer by stanbon(75887)

gina has $2 more than three times the amount fred has. when gina gave fred $5 fred then had one half asa much as gina. how much did gina have at first
--------
Equations:
g = 3f + 2
f+5 = (1/2)(g-5)
-----
Rearrange:
g - 3f = 2
g - 2f = 15
-------
Subtract and solve for "f":
f = $13 (amt. fred started with)
g = 3f+2 = $41 (amt. gina started with)
----------
Cheers,
Stan H.
-----------